### v1

# Modified Lepage-type test statistics for the weak null hypothesis

Abstract
Detecting simultaneous shifts in the location and scale of two populations is a challenging problem in statistical research. A common way to address this issue is by combining location and scale test statistics. A well-known example is the Lepage test, which combines the Wilcoxon-Mann-Whitney test for location with the Ansari-Bradley test for scale. However, the Wilcoxon-Mann-Whitney test assumes that the population variances are equal, while the Ansari-Bradley test assumes the population medians are equal. This study introduces new approaches that combine recent methodological advances to relax these assumptions. We incorporate the Fligner-Policello test, a distribution-free alternative to the Wilcoxon-Mann-Whitney test that does not require the assumption of equal variances. The Fligner-Policello test is further enhanced by the Fong-Huang method, which provides an improved variance estimation. Additionally, we propose a new variance estimator for the Ansari-Bradley test, thereby eliminating the need for the equal medians assumption. These methodological modifications are integrated into the Lepage framework to operate under a weak null hypothesis. Simulation results suggest that these new tests are promising candidates for location-scale testing. The practical utility of the proposed tests is then demonstrated through an analysis of four real-world biomedical datasets. These empirical applications confirm the robustness and reliability of the modified tests for the two-sample independent location-scale problem.
Keywords: Fligner-Policello estimator; Fong-Huang estimator; Lepage-type statistics; Location-scale model; power of test; variance estimators. 11footnotetext: corresponding author: abid0100@gmail.com
1 Introduction
The comparison of treatment and control groups is a common practice in biomedical research and related disciplines. Standard two-sample tests, such as the Student’s tt–test and FF–ratio test, are frequently employed to evaluate the effects of different treatments, such as varying medication doses, on two independent groups. When the populations are normally distributed, these tests are considered the uniformly most powerful unbiased tests. However, the underlying assumption of normality for these tests is often not met in practice. Consequently, nonparametric two-sample tests are recommended, particularly in situations involving small sample sizes, where the verification of distributional assumptions is challenging or infeasible. These tests are preferred due to their lack of distributional assumptions within each sample and their demonstrated robustness under diverse conditions, as discussed by Blair and Higgins (1980) and Van der Vaart (2000). Among the numerous available statistical tests for comparing the locations of two independent samples, the nonparametric Wilcoxon-Mann-Whitney UU–test is frequently applied in a variety of settings (Letshedi et al. 2021; Lin et al. 2021 and Dao 2022). Ansari-Bradley test is a nonparametric test that is commonly used to compare the dispersion of two independent samples. It has become one of the core statistically analytical tools in most research and practical applications (Mukherjee and Marozzi 2019; Murakami and Neuhäuser 2025 and Hussain 2025).
Simultaneous variation in location and scale is a common occurrence in statistical analysis (see e.g., Neuhäuser 2002 and Murakami and Neuhäuser 2025). For instance, in randomized controlled trials, the null hypothesis of no treatment effect often assumes homogeneous variances among the different treatment groups. However, the presence of heteroscedasticity can itself be an indicator of an underlying treatment effect (Neuhäuser, 2002). A compelling example can be found in a clinical trial analyzed by Neuhäuser (2001), which investigated the frequency of exacerbations in patients with chronic obstructive pulmonary disease (COPD). In this context, a treatment effect could manifest a shift in both the location and scale parameters. This observation has spurred significant research into developing distribution-free tests capable of simultaneously detecting differences in both location and scale between two populations. A prominent example of such a test is the sum of the squared standardized Wilcoxon-Mann-Whitney UU–test and Ansari-Bradley CC–test statistics, a method introduced by Lepage (1971).
Statistical testing typically focuses on the strong null hypothesis, which posits an exact equality between populations. However, a weak null hypothesis, which concerns a relative marginal effect, can be relevant in certain situations (Brunner et al. 2002 and Fong and Huang 2019). The Wilcoxon-Mann-Whitney UU–test, for instance, cannot directly test this weak null hypothesis because the sampling distribution of its test statistic is not distribution-free under this condition (Chung and Romano 2013 and Chung and Romano 2016). To address this limitation, Fligner and Policello (1981) developed an alternative to the UU–test. They introduced a data-based approximation for the variance of the ranks. Their resulting test, known as the Fligner-Policello (FP) test, is calculated using the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$. Here, $\widehat{Var}_{FP}(U)$ represents a consistent variance estimator, which will be described in detail in Section 2. This test statistic can be approximated by a standard normal distribution or by a permutation distribution. When an additional distributional symmetry condition is met for both random variables $X$ and $Y$, the weak null hypothesis becomes equivalent to testing for the equality of population medians. Furthermore, the FP test is generally preferred over the UU–test when testing the weak null hypothesis without the assumption of homogeneous variances (e.g., Siegel and Castellan 1988; Zumbo and Coulombe 1997 and Mickelson 2013). The FP test, therefore, provides a robust alternative for comparing population locations under less restrictive assumptions.
In this paper, we propose several modifications to the Lepage test statistic to improve its power for the two-sample location-scale problem. These changes are specifically designed to maintain or increase statistical power under weak null hypotheses-that is, when unequal variances exist in location-testing or unequal medians exist in scale-testing. One of our approaches introduces the FP test as an alternative to the traditional Wilcoxon-Mann-Whitney for assessing the location component of the Lepage test. Further, we incorporate two key methodological advancements in variance estimation under less restrictive assumptions. First, we utilize the Fong-Huang (FH) estimator, which corrects a known error in the original FP estimator (Fong and Huang, 2019). This FH test enhances the accuracy to detect the location-difference, allowing it to be reliably compared to standard normal approximation as $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$. Second, we propose a novel empirical variance estimator for the Ansari-Bradley $C$ statistic. It can be implemented through standard normal approximation as $\big(C-E_{0}(C)\big)/\sqrt{\widehat{Var}(C)}$, eliminates the need for the assumption of equal medians. The Lepage-type test statistics developed in this study are constructed by incorporating the variance estimators of the $U$ and $C$ statistics. The subsequent sections of this paper provide the specific details and derivations of these new test statistics.
This study is structured as follows: In Section 2, we begin with the details of the $U$ statistic, its variance, and providing an overview of existing estimators proposed by Fligner and Policello (1981) and Fong and Huang (2019). Section 3 introduces the $C$ statistic, its variance, and a newly proposed variance estimator. Section 4 then presents the location-scale model and the Lepage-type statistics that form the basis of our analysis. Section 5 employs a Monte Carlo simulation to evaluate the performance of our proposed tests with classical Lepage test. In Section 6, we demonstrate the practical application of these tests using four real-world datasets. Finally, Section 7 summarizes the key findings of the research along with some future directions.
2 The $U$ statistic
2.1 The classical $U$ statistic
The Wilcoxon-Mann-Whitney UU–test statistic is based on two independent random samples: $X=(X_{1},X_{2},\ldots,X_{m})$ (independent and identically distributed (i.i.d.) observations from a continuous population) and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ (i.i.d. observations from another continuous population). Let the distribution functions of $X$ and $Y$ be denoted by $\mathbb{F}$ and $\mathbb{G}$, respectively. The $U$ statistic can be computed using either the sum of ranks of the $Y$ observations in the combined sample or a $U$ statistic
$U=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}I\big(X_{i}<Y_{j}\big).$ (1)
A trade-off exists between computational efficiency and clarity of statistical properties when considering the rank sum and $U$ statistic formulations. These two forms are theoretically equivalent, differing only by constant factors (see, for example, Hollander et al. 2014). While the rank sum formulation is computationally advantageous, expressing the statistic as a $U$ statistic, as in Eq. (1), highlights its consistency and minimum variance unbiasedness in estimating $P(Y>X)$ (Lehmann, 1951).
2.2 Variance of $U$ statistic
The $U$ statistic is a nonparametric test, therefore it does not rely on specific distributional assumptions for the independent variables $X$ and $Y$. Under the null hypothesis ($X$ and $Y$ are identically distributed), the distribution of the $U$ statistic is solely determined by the sample sizes $m$ and $n$. Specifically, $E_{0}(U)=1/2$ and the variance of $U$ statistic is
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}Var\bigg(\sum_{i=1}^{m}\sum_{j=1}^{n}I\big(X_{i}<Y_{j}\big)\bigg)$
$=$ $\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{i^{*}=1}^{m}\sum_{j^{*}=1}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i^{*}}<Y_{j^{*}}\big)\bigg).$
The indices $i$ and $j$, and $i^{*}$ and $j^{*}$ can relate in four distinct ways. When $i\neq i^{*}$ and $j\neq j^{*}$, the indicator functions $I(X_{i}<Y_{j})$ and $I(X_{i^{*}}<Y_{j^{*}})$ are independent. Consequently, the total variance can be decomposed into three additive components
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j\neq j^{*}}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i}<Y_{j^{*}}\big)\bigg)$ (2)
$+\frac{1}{(mn)^{2}}\sum_{i\neq i^{*}}^{m}\sum_{j=1}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i^{*}}<Y_{j}\big)\bigg)$
$+\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j=1}^{n}Var\bigg(I\big(X_{i}<Y_{j}\bigg).$
This is made explicit by noting that
$Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i}<Y_{j^{*}}\big)\bigg)$ $=$ $E\bigg(I\big(X_{i}<Y_{j}\big)I\big(X_{i}<Y_{j^{*}}\big)\bigg)-E\bigg(I\big(X_{i}<Y_{j}\big)\bigg)E\bigg(I\big(X_{i}<Y_{j^{*}}\big)\bigg)$
$=$ $E\bigg(E\big(I(X_{i}<Y_{j})I(X_{i}<Y_{j^{*}})|X_{i}\big)\bigg)-E\bigg(E\big(I(X_{i}<Y_{j})|X_{i}\big)\bigg)^{2}$
$=$ $E\bigg(S_{Y}^{2}(X_{i})\bigg)-E^{2}\bigg(S_{Y}(X_{i})\bigg),$
where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus we have
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}\bigg(m(n^{2}-n)Var_{0}\big(S_{Y}(X_{i})\big)+n(m^{2}-m)Var_{0}\big(S_{X}(Y_{i})\big)+mnVar_{0}\big(I(X_{i}<Y_{j})\big)\bigg)$ (3)
$=$ $\frac{1}{(mn)^{2}}\bigg(m(n^{2}-n)Var_{0}\big(F_{Y}(X_{i})\big)+n(m^{2}-m)Var_{0}\big(F_{X}(Y_{i})\big)+mnVar_{0}\big(I(X_{i}<Y_{j})\big)\bigg)$
$=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{m}Var_{0}\big(F_{Y}(X_{i})\big)+\bigg(1-\frac{1}{m}\bigg)\frac{1}{n}Var_{0}\big(F_{X}(Y_{i})\big)+\frac{1}{mn}Var_{0}\big(I(X_{i}<Y_{j})\big),$
where $S_{X}(Y_{i})=E_{X|Y_{i}}\big(I(Y_{i}<X)\big)$. Under the strong null hypothesis, the cumulative distribution functions $F_{Y}(X_{i})$ and $F_{X}(Y_{i})$ are uniformly distributed random variables. Consequently, the expression evaluates to
$Var_{0}(U)$ $=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{12m}+\bigg(1-\frac{1}{m}\bigg)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{n-1}{12mn}+\frac{m-1}{12mn}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\bigg).$
For sufficiently large $m$ and $n$, the standardized statistic $(U-1/2)/\sqrt{Var_{0}(U)}$ asymptotically follows the standard normal (Hollander et al. 2014), enabling approximate pp–value calculations.
2.3 Fligner and Policello estimator
Several procedures exist for testing the equality of two medians when the populations exhibit differing shapes. The method proposed by Fligner and Policello (1981) is particularly useful as it requires fewer assumptions regarding the distributional forms of the two populations. Their estimator for the variance of the $U$ statistic can be expressed as (Fong and Huang 2019)
$\widehat{Var}_{FP}(U)$ $=$ $\frac{(m-1)n^{2}}{(mn)^{2}}\widehat{Var}(\hat{F}_{Y}(X_{i}))+\frac{(n-1)m^{2}}{(mn)^{2}}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{mn}{(mn)^{2}}\widehat{Var}\big(I(X_{i}<Y_{j})\big)$
$=$ $\bigg(1-\frac{1}{m}\bigg)\frac{1}{m}\widehat{Var}(\hat{F}_{Y}(X_{i}))+\bigg(1-\frac{1}{n}\bigg)\frac{1}{n}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{1}{mn}P_{m}\big(\hat{F}_{Y}(X_{i})\big)P_{n}\big(\hat{F}_{X}(Y_{i})\big).$
In the above expression $\widehat{Var}$ is denoted as the estimator of $Var_{0}$. Furthermore, $\widehat{Var}\big(I(X_{i}<Y_{j})\big)$ is given by $P_{m}\big(\hat{F}_{Y}(X_{i})\big)P_{n}\big(\hat{F}_{X}(Y_{i})\big)$, where $P_{m}$ represents the empirical average operator with respect to the $m$ observations $X_{1},X_{2},\cdots,X_{m}$, and $P_{n}$ represents the empirical average operator with respect to the $n$ observations $Y_{1},Y_{2},\cdots,Y_{n}$. In the case where $m\neq n$, this estimator yields the following estimate under the strong null hypothesis.
$\widehat{Var}_{FP}(U)$ $=$ $\big(1-\frac{1}{m}\big)\frac{1}{12m}+\big(1-\frac{1}{n}\big)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}+\frac{1}{n}-\frac{1}{m^{2}}-\frac{1}{n^{2}}+\frac{3}{mn}\bigg),$
which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$. However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$.
2.4 Fong and Huang estimator
Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (3), defined as
$\widehat{Var}_{FH}(U)$ $=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{m}\widehat{Var}\big(\hat{F}_{Y}(X_{i})\big)+\bigg(1-\frac{1}{m}\bigg)\frac{1}{n}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{1}{mn}\widehat{Var}\big(I(X_{i}<Y_{j})\big).$
Assuming the strong null hypothesis, the following estimate is obtained
$\widehat{Var}_{FH}(U)$ $=$ $\big(1-\frac{1}{n}\big)\frac{1}{12m}+\big(1-\frac{1}{m}\big)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}-\frac{1}{mn}+\frac{1}{n}-\frac{1}{mn}+\frac{3}{mn}\bigg),$
which is equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$, when $m\neq n$.
3 The $C$ statistic
3.1 The classical $C$ statistic
The $C$ statistic, proposed by Ansari and Bradley (1960), is used to assess a two-sample dispersion problem. Its application requires the two populations to have identical medians. To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},~Z_{2},~\cdots,~Z_{N}\}$. Ranks are then assigned symmetrically from the tails of the ordered sample. Rank 1 is assigned to the minimum ($Z_{1}$) and maximum ($Z_{N}$) values, rank 2 to the second smallest ($Z_{2}$) and second largest ($Z_{N-1}$), and so on. The array of assigned ranks, denoted as $S_{N}$, depends on whether the combined sample size $N$ is even or odd. If $N$ is an even integer, the ranks are given by
$S_{N}=\bigg\{1,~2,~3,~\cdots,~\frac{N}{2},~\frac{N}{2},~\cdots,~3,~2,~1\bigg\}.$
If $N$ is an odd integer, the ranks are given by
$S_{N}=\bigg\{1,~2,~3,~\cdots,\frac{N-1}{2},~\frac{N+1}{2},~\frac{N-1}{2},~\cdots,~3,~2,~1\bigg\}.$
Let $R_{i}$ be the rank assigned to the $i$th value from the $Y$ sample. The test statistic, $C$, is defined as the sum of the ranks assigned to all values in the $Y$ sample
$C=\sum_{i=1}^{n}R_{i},$ (4)
where $n$ is the size of the $Y$ sample. An alternative formulation, equivalent to Eq. (4) and more useful for mathematical considerations, is given by
$C=\sum_{i=1}^{k}i\lambda_{i}+\sum_{i=k+1}^{N}(N+1-i)\lambda_{i},$
where
$\lambda_{i}=\begin{cases}1;&\ \ \text{if $Z_{i}\in$ $Y$ }\\ &\\ 0;&\ \ \text{otherwise,}\end{cases}$
and $k$ is the largest integer in $\frac{N+1}{2}$. Small values of $C$ suggest a larger dispersion for the $Y$ sample, whereas large values of $C$ indicate a larger dispersion for the $X$ sample.
3.2 Variance of $C$ statistic
Under the null hypothesis ($H_{0}$), all $\binom{N}{n}$ possible assignments of ranks to the combined sample are equally probable. For large sample approximation, the statistic $\frac{C}{n}$ represents the average of the ranks assigned to the $Y$ sample. Its distribution is identical to the distribution of the sample mean for a random sample of size $n$ drawn without replacement from the finite population of ranks. For an even $N$, the expected value of the statistic, $C$, is given by
$E_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\bigg(\frac{N}{2}+1\bigg)}{2}=\frac{N+2}{4}$
$E_{0}(C)$ $=$ $\frac{n(N+2)}{4}.$
The variance of $C$ can be calculated as
$Var_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}Cov(R_{i},R_{j})$
$=$ $\frac{1}{n^{2}}\sum_{i=1}^{n}Var(R_{i})+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j\neq i}Cov(R_{i},R_{j})$
$=$ $\frac{\sigma^{2}}{n}+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j\neq i}\bigg(\frac{-\sigma^{2}}{N-1}\bigg)=\frac{\sigma^{2}}{n}-\frac{n(n-1)}{n^{2}}\bigg(\frac{\sigma^{2}}{N-1}\bigg)$
$Var_{0}(C)$ $=$ $\sigma^{2}\bigg(\frac{n(N-n)}{N-1}\bigg),$ (5)
where
$\sigma^{2}$ $=$ $\frac{2}{N}\sum_{i=1}^{N/2}i^{2}-\mu^{2}=\frac{N^{2}-4}{48}.$
Therefore, the final expression for the variance is
$Var_{0}(C)$ $=$ $\bigg(\frac{N^{2}-4}{48}\bigg)\bigg(\frac{n(N-n)}{N-1}\bigg)=\frac{mn(N^{2}-4)}{48(N-1)}.$ (6)
For an odd $N$, the expected value and variance can easily be calculated as
$E_{0}(C)$ $=$ $\frac{n(N+1)^{2}}{4N}$
$Var_{0}(C)$ $=$ $\frac{mn(N+1)\big(N^{2}+3\big)}{48N^{2}}.$ (7)
These results are consistent with standard theory for the mean of a sample from a finite population (see, for example, Hollander et al. 2014).
For any NN (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (see, for example, Ansari and Bradley 1960 and Randles and Wolfe 1979).
3.3 A suggested variance estimator
This section presents a procedure for testing the dispersion of two independent samples, particularly when the underlying assumption regarding populations medians is violate. A key component of this method involves estimating the population variance, $\sigma^{2}$, using results from a sample. The expected value of this estimator is given by
$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\times\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
Using the known expression for the variance of $C$ (3.2) and the estimator $\hat{\sigma}^{2}$, the variance of $C$ can be estimated as follows
$\widehat{Var}(C)=\hat{\sigma}^{2}\times\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\times\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\times\bigg(\frac{n^{2}(N-n)}{N(n-1)}\bigg),$ (8)
where $\hat{\sigma}^{2}$ represents the observed variance of the ranks of sample $Y$. The proposed estimator is equally applicable for two different variances, as described in Eq. (6) and Eq. (3.2), regardless of whether the sample size $N$ is even or odd.
Following the strategy from Fong and Huang (2019), a simulation study is conducted to evaluate the performance of the proposed variance estimator. Data are generated from two different distributions
  1. $\bullet$
    A logistic distribution: $X\sim\mathcal{L}(0,1)$ versus $Y\sim\mathcal{L}(0,1)$,
  2. $\bullet$
    An exponential distribution: $X\sim\mathcal{E}(1)$ versus $Y\sim\mathcal{E}(1)$.
To compute the true variance of the $C$ statistic, Eq. (6) or (3.2) is used, depending on whether the sample size ($N$) is even or odd. This true variance is then compared to the average of the variance estimates obtained using Eq. (8). Table 1, based on 10,000 replications, shows that the proposed estimator provides accurate and consistent estimates of the true variance for both probability distributions. This consistency suggests that the proposed estimator is a consistent and distribution-free variance estimator of $C$ statistic. The test statistic for the weak null hypothesis is therefore given by
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
Table 1: Comparison of $C$ statistic variance and its estimator.
                                                                  Sample sizes
$m=n=5$ $m=n=10$ $m=20,n=15$ $m=30,n=25$ $m=n=30$ $m=n=50$
$Var_{0}(C)$ 5.5556 43.4211 224.4485 874.1319 1142.7970 5258.8380
Logistic distribution
$\widehat{Var}(C)$ 5.5671 43.4862 224.8931 874.9312 1142.5058 5257.0210
Exponential distribution
$\widehat{Var}(C)$ 5.5507 43.5181 225.4013 873.8417 1143.1610 5256.8125
Using a permutation-based approach, we calculated the exact critical values for the $C^{*}$ and $C_{P}^{*}$ test statistics. Our results indicated that both statistics have identical critical values across various sample size combinations. Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$. The observed similarity in critical values suggests that the asymptotic normality of the $C_{P}^{*}$ test statistic under the null hypothesis is consistent with that of the $C^{*}$ test statistic.
3.4 Foundation in Slutsky’s theorem
Slutsky’s theorem is a foundational principle in probability theory that is instrumental in establishing the asymptotic normality of a statistic, particularly when it incorporates a consistent estimator. The theorem states that if a sequence of random variables, $A_{n}$, converges in distribution, $D$, and another sequence of random variables, $B_{n}$, converges in probability to a constant, $a$, then the product of these sequences, $A_{n}B_{n}$, converges in distribution to $aD$ (see e.g., Casella and Berger 2002, p. 239–240). This can be expressed as
$\text{If}~~A_{n}\overset{d}{\longrightarrow}D~~\text{and}~~B_{n}\overset{p}{\longrightarrow}a,~~\text{then}~~A_{n}B_{n}\overset{d}{\longrightarrow}aD.$
We can apply this theorem to demonstrate the asymptotic normality of the test statistic, $C_{P}^{*}$. This statistic can be expressed as the product of two components in the following form
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\times\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
The first component of Eq. (9), $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal distribution, whereas the second component, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1. According to Slutsky’s theorem, their product, the test statistic $C_{P}^{*}$ is asymptotically normally distributed under a weak $H_{0}$
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
4 Location-scale models
4.1 Hypotheses
Let $X=(X_{1},X_{2},\ldots,X_{m})$ and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ denote two i.i.d. random samples of sizes $m$ and $n$, respectively. Let $\mathbb{F}$ and $\mathbb{G}$ represent the continuous distribution functions associated with the samples $X$ and $Y$, respectively. The null hypothesis under consideration posits that both samples are drawn from the same population as
$H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta),~~~\forall~\theta\in(-\infty,\infty).$
Our aim is to assess the null hypothesis against the location-scale alternative
$H_{1}:\mathbb{F}(\theta)=\mathbb{G}\left(\frac{\theta+\Delta}{\tau}\right),$
where
$\Delta\neq 0,~~~\tau\neq 1,~~~\Delta\in(-\infty,\infty),~~~\tau\in(0,\infty).$
In hypothesis $H_{1}$, $\Delta$ represents the location difference, while $\tau$ represents the scale difference between the distributions $\mathbb{F}$ and $\mathbb{G}$.
4.2 The Lepage-type test statistics
The classical Lepage statistic, introduced by Lepage (1971), combines the Wilcoxon-Mann-Whitney UU–test for location difference and the Ansari-Bradley CC–test for scale variation. The $U$ statistic, defined in Eq. (1), is applicable to the location model $\mathbb{F}(\theta)=\mathbb{G}(\theta+\Delta)$, as formalized by the following hypotheses
$H_{0}:\Delta=0~~~~~~\text{versus}~~~~~~H_{1}:\Delta<0,~\Delta>0~~\text{or}~~\Delta\neq 0.$
The CC–test, within the framework of a location-scale model, relates to the scale model $\mathbb{F}(\theta)=\mathbb{G}(\theta/\tau)$ by imposing the restriction that the location parameter $\Delta$ equals zero. This restriction gives the subsequent hypotheses
$H_{0}:\tau=1~~~~~~\text{versus}~~~~~~H_{1}:\tau<1,~\tau>1~~\text{or}~~\tau\neq 1.$
Then Lepage statistic is defined as
$L_{0}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}.$
Lepage demonstrated that the statistics $U$ and $C$ are linearly uncorrelated inside the $L_{0}$ statistic framework for all $m$ and $n$; hence, this test under $H_{0}$ asymptotically conforms to a $\chi^{2}$ distribution with 2 degrees of freedom (Lepage, 1971).
This research presents a methodology for constructing five Lepage-type test statistics to assess a weak null hypothesis. These statistics are constructed using the two estimators for $Var_{0}(U)$ detailed in Section 2 and the single estimator for $Var_{0}(C)$ described in Section 3. These estimators are formulated within the framework of the Lepage test statistic as follows
$L_{1}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{2}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{3}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
$L_{4}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
and
$L_{5}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}.$
Under the null hypothesis ($H_{0}$), the test statistic $L_{i}$ (for all $i=1,2,\cdots,5$) is assumed to follow an asymptotic chi-squared distribution with two degrees of freedom. This is because the constituent statistics, $U$ and $C$, are independent of each other and each of them asymptotically follows a standard normal distribution, as discussed in Sections 2 and 3.
4.3 Critical values for small sample sizes
The Lepage-type test statistics, when used with a normal approximation, are expected to maintain the nominal Type I error rate under the null hypothesis, provided that the sample sizes are sufficiently large. However, for small sample sizes, a permutation-based methodology offers a more robust approach for controlling the Type I error rate. To construct the permutation distribution, an ideal approach would be to enumerate all possible ways of reassigning group labels among the $N$ total samples and calculate the test statistic for each permutation. Due to the computational infeasibility of exhaustive enumeration for even moderately sized samples, a Monte Carlo approach is typically employed. This method involves generating a large number of random permutations to approximate the null distribution.
In this study, we utilized the R package NSM3, which is specifically designed to provide exact and Monte Carlo permutation-based methods for various nonparametric statistical tests, including the Lepage test. The function $cLepage(\alpha,m,n)$” is used to compute the critical values. For all modified Lepage test statistics, we conducted a simulation study under small-sample scenario, with 100,000 simulations. To ensure the consistency of the results, we varied the underlying probability distributions under the null hypothesis, including the normal, uniform, logistic, and Laplace distributions. Across all distributions, the resulting critical values remained consistent. These values are presented in Table 2.
Table 2: Critical values for right-tailed Lepage-type test statistics at a 5% significance level.
                                                       Test statistic
$m$ $n$ $L_{0}$ $L_{1}$ $L_{2}$ $L_{3}$ $L_{4}$ $L_{5}$
5 5 5.3345 8.8948 8.8948 7.2012 10.3906 10.3906
6 5 5.5269 7.7793 7.9803 7.7727 12.4460 12.4467
6 6 5.7692 6.8571 6.8571 6.8173 11.2084 11.2084
7 5 5.5720 7.9068 7.9068 8.5803 11.5886 11.5886
7 7 5.6541 6.8855 6.8855 7.0367 9.0802 9.0802
8 5 5.5037 7.5440 7.4924 8.6301 12.2734 12.2084
8 8 5.6775 6.6280 6.6280 6.8773 8.3711 8.3711
9 5 5.4444 7.2574 7.2574 8.3956 12.0756 12.0756
10 5 5.4468 7.3250 7.3355 9.0133 11.6773 11.6655
10 10 5.7436 6.5368 6.5719 6.5545 7.5905 7.5905
5 Size and power comparison
A Monte Carlo simulation study is conducted in R (version 4.5.1) to assess the performance of Lepage-type test statistics. Power levels are estimated based on 10,000 simulations, with the significance level for rejecting the null hypothesis ($H_{0}$) set at $0.05$. The power performances of the proposed $L_{i}$ (for all $i=1,2,\cdots,5$) test statistics are compared to that of the classical $L_{0}$ test under simultaneous location-scale shift. For the sake of brevity, the significance level and estimated power of $L_{0}$ and $L_{i}$ (for all $i=1,2,\cdots,5$) are examined for two sample size combinations, $(m,n)$: $(10,10)$ and $(40,30)$. The hypotheses under consideration are $H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta)$ versus $H_{1}:\mathbb{F}(\theta)=\mathbb{G}(\frac{\theta-\Delta}{\tau})$, where $\Delta\neq 0$ and $\tau\neq 1$. To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ are used to represent the following distributions.
  1. $\bullet$
    $\mathbb{F}\sim\mathcal{E}(0.5)$ versus $\mathbb{G}\sim\mathcal{E}(\theta)$: The exponential distribution is a continuous probability distribution characterized by its right-skewed nature. It is frequently employed to model the duration of time until a specific event occurs, such as the time until a mechanical component experiences failure, the inter-arrival time of customers at a service point, or the decay time of a radioactive particle.
  2. $\bullet$
    $\mathbb{F}\sim\chi^{2}_{(1)}$ versus $\mathbb{G}\sim\chi^{2}_{(\theta)}$: The chi-square distribution is a continuous, asymmetric probability distribution characterized by a rightward skew. It is a fundamental tool in inferential statistics, particularly for applications in hypothesis testing.
  3. $\bullet$
    $\mathbb{F}\sim\mathcal{G}(1,1)$ versus $\mathbb{G}\sim\mathcal{G}(1,\theta)$: The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew. It is widely employed to model waiting times and other non-negative random variables, finding extensive application in diverse fields, such as reliability engineering, queueing theory, insurance and hydrology, etc.
  4. $\bullet$
    $\mathbb{F}\sim\mathcal{B}(2,2)$ versus $\mathbb{G}\sim\mathcal{B}(2,\theta)$: The beta distribution is a family of continuous probability distributions defined on the interval $[0,1]$. It is primarily utilized to model the behavior of random variables representing proportions, percentages, or probabilities. Unlike unimodal distributions like the normal or Laplace distributions, the beta distribution exhibits remarkable flexibility in its shape. Its form can vary significantly, ranging from a straight line to U-shaped, J-shaped, or even bell-shaped, depending on the specific values of its two parameters. This adaptability makes it a valuable tool in various fields. The beta distribution is well-suited for several applications, including Bayesian inference, project management and reliability engineering.
  5. $\bullet$
    $\mathbb{F}\sim\mathcal{U}(0,1)$ versus $\mathbb{G}\sim\mathcal{U}(0,\theta)$: A continuous uniform distribution is a probability distribution where every outcome within a defined interval has an equal probability of occurrence. This core property makes it a valuable baseline model in statistical analysis and simulations. The uniform distribution is widely used in situations where there is no a priori reason to believe that any single outcome is more likely than another. It is widely employed for statistical analysis, simulations, and the modeling of real-world phenomena.
Simulation results for the exponential distribution are presented in Table 3, examining both small $(m=n=10)$ and large $(m=40,n=30)$ sample sizes. The first column, which corresponds to the case of identical distributions, shows the simulated Type I error rate. A size correct test must have a Type I error rate close to the nominal significance level of $0.05$. For large samples, the $L_{0}$ test appears slightly conservative, while the proposed tests are slightly liberal. This is likely due to the use of asymptotic critical values rather than permutation-based ones. All other columns demonstrate that the proposed tests have significantly higher power than $L_{0}$, with the exception of the $L_{3}$ test. These results suggest that the proposed Lepage-type tests are a suitable alternative for analyzing exponentially distributed data.
Simulation results for the chi-square distribution are displayed in Table 4. For small samples, the Type I error rates are close to the nominal level. However, for large samples, a slightly liberal behavior is observed in the $L_{3}$, $L_{4}$ and $L_{5}$ test statistics. The power comparison shows that the proposed tests consistently maintain higher power across all findings for both small and large sample sizes. As observed previously with the exponential distribution, the proposed $L_{3}$ test statistic demonstrates the lowest power.
An analysis of the Gamma distribution is presented in Table 5. The simulation study indicates that the proposed statistics consistently outperform the classic $L_{0}$ statistic for both small and large sample comparisons, with the exception of $L_{3}$. For Type I error, all tests maintain the rate near $0.05$, though $L_{4}$ and $L_{5}$ exhibit a slightly liberal behavior in the large sample comparison. In summary, the proposed test statistics generally demonstrate greater efficiency than the $L_{0}$ statistic for estimating parameters of the Gamma distribution under both small and large sample conditions.
Table 3: Simulated actual Type I error rate and power of $\mathcal{E}(0.5)$ versus $\mathcal{E}(\theta)$.
$\theta$
$m$ $n$ Test 0.5 1.0 1.5 2.0 2.5
10 10 $L_{0}$ 0.0526 0.1891 0.4071 0.5992 0.7343
$L_{1}$ 0.0509 0.2169 0.4758 0.6680 0.7829
$L_{2}$ 0.0473 0.2127 0.4665 0.6582 0.7801
$L_{3}$ 0.0544 0.1630 0.3557 0.5233 0.6684
$L_{4}$ 0.0525 0.1958 0.4295 0.6168 0.7467
$L_{5}$ 0.0513 0.1944 0.4283 0.6217 0.7480
40 30 $L_{0}$ 0.0458 0.6164 0.9587 0.9984 0.9999
$L_{1}$ 0.0544 0.6609 0.9704 0.9987 1.0000
$L_{2}$ 0.0553 0.6568 0.9683 0.9982 1.0000
$L_{3}$ 0.0583 0.6317 0.9603 0.9970 0.9998
$L_{4}$ 0.0608 0.6615 0.9699 0.9980 0.9999
$L_{5}$ 0.0612 0.6676 0.9696 0.9975 0.9998
Table 4: Simulated actual Type I error rate and power of $\chi^{2}_{(1)}$ versus $\chi^{2}_{(\theta)}$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0534 0.1051 0.2626 0.4508 0.6387
$L_{1}$ 0.0492 0.1279 0.3203 0.5284 0.7110
$L_{2}$ 0.0496 0.1373 0.3124 0.5320 0.7165
$L_{3}$ 0.0503 0.1022 0.2135 0.3814 0.5606
$L_{4}$ 0.0536 0.1188 0.2726 0.4802 0.6525
$L_{5}$ 0.0511 0.1146 0.2831 0.4747 0.6642
40 30 $L_{0}$ 0.0480 0.3230 0.7984 0.9738 0.9978
$L_{1}$ 0.0535 0.3582 0.8320 0.9811 0.9986
$L_{2}$ 0.0521 0.3492 0.8265 0.9806 0.9987
$L_{3}$ 0.0584 0.3348 0.7925 0.9772 0.9985
$L_{4}$ 0.0604 0.3709 0.8323 0.9815 0.9987
$L_{5}$ 0.0583 0.3659 0.8309 0.9790 0.9984
Table 5: Simulated actual Type I error rate and power of $\mathcal{G}(1,1)$ versus $\mathcal{G}(1,\theta)$.
$\theta$
$m$ $n$ Test 1.0 2.0 3.0 4.0 5.0
10 10 $L_{0}$ 0.0496 0.1882 0.4070 0.6077 0.7246
$L_{1}$ 0.0490 0.2236 0.4787 0.6627 0.7916
$L_{2}$ 0.0475 0.2220 0.4770 0.6621 0.7812
$L_{3}$ 0.0504 0.1638 0.3523 0.5333 0.6714
$L_{4}$ 0.0533 0.1951 0.4282 0.6200 0.7492
$L_{5}$ 0.0478 0.1962 0.4241 0.6216 0.7501
40 30 $L_{0}$ 0.0496 0.6163 0.9621 0.9971 0.9997
$L_{1}$ 0.0500 0.6651 0.9712 0.9985 1.0000
$L_{2}$ 0.0494 0.6504 0.9677 0.9989 1.0000
$L_{3}$ 0.0536 0.6255 0.9633 0.9975 0.9999
$L_{4}$ 0.0607 0.6713 0.9716 0.9982 0.9999
$L_{5}$ 0.0612 0.6735 0.9731 0.9979 0.9999
In a simulation study of the beta distribution, the proposed test statistics clearly outperform the traditional $L_{0}$ statistic in detecting simultaneous shifts under both small and large sample comparisons, as shown in Table 6. The $L_{4}$ and $L_{5}$ tests are slightly liberal in the large sample comparison. The proposed tests show superior power compared to the classical Lepage test. For this distribution, $L_{3}$ has the lowest power in the small sample investigation but performs similarly to $L_{0}$ in the large sample case. These findings suggest that the proposed tests offer a robust alternative to $L_{0}$ for analyzing parameters of the beta distribution.
For the uniform distribution, all competing tests show similar performance in both sample scenarios, see Table 7. Consistent with previous findings, the proposed $L_{3}$, $L_{4}$, and $L_{5}$ tests are slightly liberal for large sample investigation.
Table 6: Simulated actual Type I error rate and power of $\mathcal{B}(2,2)$ versus $\mathcal{B}(2,\theta)$.
$\theta$
$m$ $n$ Test 2.0 3.0 4.0 5.0 6.0
10 10 $L_{0}$ 0.0495 0.1221 0.2927 0.4647 0.6441
$L_{1}$ 0.0512 0.1414 0.3333 0.5486 0.7104
$L_{2}$ 0.0504 0.1391 0.3413 0.5482 0.7093
$L_{3}$ 0.0487 0.1033 0.2407 0.3971 0.5708
$L_{4}$ 0.0518 0.1252 0.3074 0.4871 0.6692
$L_{5}$ 0.0503 0.1204 0.3059 0.4910 0.6614
40 30 $L_{0}$ 0.0509 0.3688 0.8383 0.9843 0.9989
$L_{1}$ 0.0512 0.4127 0.8698 0.9874 0.9981
$L_{2}$ 0.0511 0.4020 0.8659 0.9852 0.9990
$L_{3}$ 0.0525 0.3737 0.8420 0.9837 0.9988
$L_{4}$ 0.0633 0.4142 0.8763 0.9855 0.9992
$L_{5}$ 0.0577 0.4188 0.8683 0.9856 0.9990
Table 7: Simulated actual Type I error rate and power of $\mathcal{U}(0,1)$ versus $\mathcal{U}(0,\theta)$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0493 0.2565 0.5619 0.7444 0.8454
$L_{1}$ 0.0491 0.2794 0.5699 0.7462 0.8498
$L_{2}$ 0.0440 0.2813 0.5627 0.7524 0.8545
$L_{3}$ 0.0488 0.2417 0.5213 0.7029 0.8130
$L_{4}$ 0.0480 0.2448 0.5402 0.7318 0.8339
$L_{5}$ 0.0549 0.2556 0.5362 0.7253 0.8359
40 30 $L_{0}$ 0.0435 0.7935 0.9888 0.9993 1.0000
$L_{1}$ 0.0489 0.8122 0.9904 0.9997 1.0000
$L_{2}$ 0.0526 0.8054 0.9912 0.9994 1.0000
$L_{3}$ 0.0572 0.7874 0.9879 0.9994 1.0000
$L_{4}$ 0.0629 0.7997 0.9897 0.9994 0.9999
$L_{5}$ 0.0638 0.7993 0.9894 0.9995 1.0000
6 The empirical study
The focus of this section is to explore the application and evaluation of proposed Lepage-type tests. We compare their performance to the traditional Lepage test using four illustrative datasets, seeking to identify significant changes in location, scale, or both. Figure 1 displays box-plots of the data, and Table 8 provides detailed dataset descriptions.
Our initial dataset, derived from Karpatkin et al. (1981) and subsequently analyzed by Hollander et al. (2014), investigates the effect of maternal steroid treatment (prednisone) on newborn platelet counts. One group of mothers received no prednisone, while the other did. The treatment aims to increase newborn platelet counts, reducing the risk of cerebral hemorrhage, by inhibiting splenic clearance of maternal antibody-coated platelets. We sought to determine if prednisone increases platelet counts. Figure 1 (upper-left panel) suggests increased variability in the treated group, prompting us to use Lepage-type tests to detect simultaneous location-scale shifts. The pp–values reported in Table 9 indicate that all competing tests detected a statistically significant difference at the 5% level of significance. This result is in agreement with the visual observations.
The second dataset is taken from Hollander et al. (2014), also investigated by Marozzi (2009) and Kössler and Mukherjee (2020). This dataset originates from a biomedical experiment designed to compare peak human plasma growth hormone levels following arginine hydrochloride infusion. Group A ($m=10$), characterized by heightened cardiovascular risk and Type A personality traits, is compared with Group B ($n=11$), exhibiting resistance to coronary disease. Visual inspection (Figure 1, upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test. The $L_{0}$, and the proposed tests, applied to this data, all yielded pp–values below 0.05 (Table 9), confirming the visual observations and aligning with prior findings (see, for example, Marozzi 2009 and Kössler and Mukherjee 2020).
The effect of thyroxine on juvenile mice growth is explored in Dataset 3 (Tasdan and Sievers, 2009). Visual comparison of the control $(m=7)$ and treatment $(n=8)$ groups’ weight measurements (daily for 30 days) in Figure 1 (bottom-left panel) reveals marked differences in distribution, including skewness, tail weight, location, and scale. These distributional characteristics suggest of a location-scale test. Therefore, $L_{0}$, and the proposed tests are employed to analyze this weight data. The resulting pp–values (Table 9) indicate statistically significant differences in weight between the control and treatment groups at the 5% significance level for all three tests.
We investigated the sensitivity of location-scale tests to detect subtle differences in cholesterol levels related to smoking status, using Dataset 4 from the Data Story Library, a dataset also analyzed by Kössler and Mukherjee (2020). This dataset compares cholesterol levels in age-matched ex-smokers (cessation within five years) and long-term smokers (25+ years). Our research question focused on whether the Lepage-type tests could identify minor variations, even when no significant differences in mean or variance are observed, as visualized in Figure 1 (bottom-right panel). As shown in Table 9, the insignificant pp–values obtained from all competing tests indicate no difference in cholesterol levels between the groups, demonstrating the effectiveness of our proposed approaches and their alignment with traditional method.
Table 8: Real life datasets.
Dataset 1: Platelet counts of newborn infants of two groups (per cubic millimeter).
Treated cases: 120 124 215 90 67 126 95 190 180 135 399 65
Control cases: 12 20 112 32 60 40 18
Dataset 2: Peak concentrations of human growth hormone in plasma following arginine
                 hydrochloride infusion.
Type-A subjects: 3.6 2.6 4.7 8.0 3.1 8.8 4.6 5.8 4.0 4.6
Type-B subjects: 16.2 17.4 8.5 15.6 5.4 9.8 14.9 16.6 15.9 5.3 10.5
Dataset 3: Thyroid data (in grams).
Control group: 0.7 1.2 1.4 2.3 1.6 0.9 1.3
Treatment group: 4.1 4.4 3.3 2.1 3.5 2.9 2.8 4.3
Dataset 4: Effect of smoking on cholesterol.
Smokers: 225 211 209 284 258 216 196 288 250 200 209 280
225 256 243 200 213 246 225 237 232 267 232 216
216 243 200 155 216 271 230 309 183 280 217 305
287 217 246 351 200 280 209
Ex-smokers: 250 134 300 249 213 310 175 174 328 160 188 321
213 257 292 200 271 227 238 163 263 192 242 249
242 267 243 217 267 218 217 183 228
Table 9: pp–values for the data examples.
                                                       Test statistic
Dataset $L_{0}$ $L_{1}$ $L_{2}$ $L_{3}$ $L_{4}$ $L_{5}$
Dataset 1 0.0030 0.0000 0.0000 0.0028 0.0000 0.0000
Dataset 2 0.0033 0.0000 0.0000 0.0033 0.0000 0.0000
Dataset 3 0.0074 0.0000 0.0000 0.0074 0.0000 0.0000
Dataset 4 0.5869 0.5879 0.5880 0.6292 0.6304 0.6304
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Box-plots of the four real datasets.
7 Concluding remarks
The two-sample location-scale problem is a frequent challenge in fields such as bioinformatics and medicine. Tests that can simultaneously detect shifts in both location and scale are often more powerful than tests that only focus on location, such as the Behrens-Fisher test. This is particularly relevant in biological research, where interventions can affect both the central tendency and the variability of a dataset, requiring methods sensitive to both parameters. Given that biological data, even after preprocessing, may deviate from a normal distribution (e.g., due to outliers, heavy tails, or skewness), nonparametric approaches are often preferred. They are more robust and efficient under these conditions due to their relaxed distributional assumptions and their suitability for small to moderate sample sizes.
Lepage (1971) test statistic is a well-established nonparametric solution for the location-scale shift problem. This study introduced several new Lepage-type statistics based on different variance estimators for the Wilcoxon-Mann-Whitney ($U$) and Ansari-Bradley ($C$) statistics, which are designed for weak null hypotheses. Simulation studies across various distributions, including uniform, right-skewed, and heavy-tailed, demonstrated that the proposed tests achieve higher power in both small and large sample investigations. The application of these new statistics to four biological datasets further manifested their practical suitability. Given their enhanced power, these proposed Lepage-type statistics are recommended for practical use.
Future work is recommended in two key areas to build upon the current findings. The first involves extending the proposed methodology to include a shape parameter, which would allow for the simultaneous evaluation of location, scale, and shape. The second area of investigation is the generalization of the proposed tests for application in multi-sample comparisons.
Disclosure statement
None declared.
Funding statement
No grant received for this research.
Ethical statement
The research work has not been previously published or submitted in any form.